# Impact of environmental interaction on bias induced circular current in a ring nanojunction

### v1

@@ III Results and discussion
almost decoupled from the backbone sites. In that case, the symmetry between the upper and lower arms of the ring is restored, resulting in a vanishing $I_{c}$.
Since the nature of circular current is greatly influenced by the interaction between the ordered region i.e., the ring system and the disordered backbone region, here it is important to check the effect of $\eta$ on $I_{c}$. In Fig. 5 we plot $I_{c}$ as a function of $W$ at three typical values of $\eta$ considering the lengthwise symmetric ring nanojunction with $N_{R}=10$. All the three curves appear identical in nature. The key observation is that, the critical $W$, where $I_{c}$ reaches it extremum, shifts towards higher $W$ with increasing $\eta$. The underlying mechanism depends on the strength of the connection between the ordered and disordered parts. A stronger connection requires a higher $W$ to decouple these two regions, from which point the symmetry between the two arms of the ring begins to reemerge.
The atypical behavior of $I_{c}$ with $W$ which is discussed above perfectly holds for the other ring-electrode
@@
junction configurations as well, which we claim from the curves shown in Fig. 6. In this figure, three junction configurations are taken into account, two are length-wise asymmetric and one for length-wise symmetric, and the results are worked out for a typical bias voltage $V=0.4\,$V. Along with these, some other configurations are also checked and in each case the dependence of $W$ on $I_{c}$ remains exactly same. It is well-known that ring-electrode configuration plays a significant role on transport behavior, due to the modification of quantum interference of the wave functions associated with different branches of the ring, and therefore, the maxima points (in the --ve side of each curve) get shifted, but the overall response of $I_{c}$ with $W$ is unchanged.
To inspect the robustness i.e., whether the above discussed nature is specific to AAH type disorder or it is general with respect to the other type of backbone disorder, here we check the behavior of $I_{c}$-$W$ curve in presence of other types of backbone disorder. We choose two other types of disorder in backbone sites: one is the Fibonacci type, and the other is the Bronze Mean (BM) one r11 ; r12 ; r13 . These are quite common examples of correlated disorders and they can be constructed using two different kinds of atomic sites, say $A$ and $B$, by arranging them according to specific rules r11 ; r12 ; r13 , unlike AAH type where all the sites are different. The inflation rule for the Fibonacci sequence is: $A\rightarrow AB$ and $B\rightarrow A$. So the first few Fibonacci generations are $A$, $AB$, $ABA$, $ABAAB$, etc. On the other hand, the inflation rule for the BM sequence is: $A\rightarrow AAAB$ and $B\rightarrow A$, and here the first few generations are: $A$, $AAAB$, $AAABAAABAAABA$, etc. For these two different atomic sites ($A$ and $B$) we refer the site energies $\epsilon_{A}$ and $\epsilon_{B}$ respectively, and their strengths are specified by the parameter $W$ (like the AAH case). Inserting the Fibonacci and BM disorders in the backbone sites, the variations of circular current with respect to $W$ are shown in Fig. 7. Here we consider a $13$-site ring, to have a comparative ring size with our previously studied AAH case ($10$-site rings cannot be obtained with these sequences). As the ring size is odd, identical arm lengths are not possible, and we couple the electrodes in such a way that the length difference between the two arms is just a single bond. For this particular figure (Fig. 7), the chosen TB parameters are also quite different compared to the other figures. It is just because to capture all the information within the varied $W$ window. We make $t_{0}$ quite large, and thus, we proportionately change other parameters such that energy band widths of the electrodes are larger than the system placed between them. In each type of backbone disorder, the variation of $I_{c}$ with $W$ is shown for three distinct bias voltages, and both for the two types of disorders the nature of $I_{c}$ with $W$ remains exactly similar with what is already obtained for the AAH case.
@@
The results analyzed so far are computed for zero temperature. For a more realistic scenario it is also crucial to examine the role of temperature. The temperature dependence enters into the current expression through the Fermi-Dirac distribution functions, $f_{S}$ and $f_{D}$, associated with the source and drain electrodes, respectively. The effect of temperature is presented in Fig. 8, where $I_{c}$-$W$ curves are shown for three distinct temperatures. The result of zero temperature is also superimposed for comparison. Apart from a slight reduction of current with temperature, we find that the overall nature of $I_{c}$-$W$ curve remains unaltered. The reduction of current due to temperature can be understood from the following arguments. At absolute zero temperature, the current is obtained by integrating the current density profile within the energy window from $E_{F}-eV/2$ to $E_{F}+eV/2$. Within this energy zone, depending on the dominating peaks and dips we get the net current. Here the chance of mutual cancellations is relatively less. On the other hand, when the temperature is finite, we need to consider the full available energy window where the chance of mutual cancellations may slightly increase though it depends on many other factors, especially $W$, and depending on the weight factor ($f_{S}-f_{D}$) we get the resultant current.
@@ IV Closing Remarks
Using a tight-binding framework to describe the quantum system, all results are derived based on wave-guide theory. The effect of backbone disorder is particularly noteworthy. As the disorder strength increases, the current magnitude initially rises, reaches a maximum, and then decreases, eventually vanishes at very high disorder strengths. These behaviors are thoroughly analyzed with appropriate mathematical results and physical explanations. Additionally, it is established that the effect of disorder remains largely unchanged regardless of the type of disorder. Temperature dependence is also discussed considering a more realistic scenario, showing that significant current can still be obtained over a wide temperature range.
Finally, we emphasize that if environmental interactions can be controlled through suitable laboratory methods, transport behavior can be selectively regulated. This potential for control could be highly interesting and important.

### v2

@@ III Results and discussion
the Fibonacci sequence is: $A\rightarrow AB$ and $B\rightarrow A$. So the first few Fibonacci generations are $A$, $AB$, $ABA$, $ABAAB$, etc. On the other hand, the inflation rule for the BM sequence is: $A\rightarrow AAAB$ and $B\rightarrow A$, and here the first few generations are: $A$, $AAAB$, $AAABAAABAAABA$, etc. For these two different atomic sites ($A$ and $B$) we refer the site energies $\epsilon_{A}$ and $\epsilon_{B}$ respectively, and their strengths are specified by the parameter $W$ (like the AAH case). Inserting the Fibonacci and BM disorders in the backbone sites, the variations of circular current with respect to $W$ are shown in Fig. 7. Here we consider a $13$-site ring, to have a comparative ring size with our previously studied AAH case ($10$-site rings cannot be obtained with these sequences). As the ring size is odd, identical arm lengths are not possible, and we couple the electrodes in such a way that the length difference between the two arms is just a single bond. For this particular figure (Fig. 7), the chosen TB parameters are also quite different compared to the other figures. It is just because to capture all the information within the varied $W$ window. We make $t_{0}$ quite large, and thus, we proportionately change other parameters such that energy band widths of the electrodes are larger than the system placed between them. In each type of backbone disorder, the variation of $I_{c}$ with $W$ is shown for three distinct bias voltages, and both for the two types of disorders the nature of $I_{c}$ with $W$ remains exactly similar with what is already obtained for the AAH case.
@@
For our chosen quantum system, it is established that electrical conduction can be monitored in different ways. In particular, backbone disorder plays a crucial role, as it causes the system to map onto an ordered-disordered separated structure, where the coupling between the ordered and disordered regions is significant. Other influential factors include the ring-electrode junction configuration, which directly affects quantum interference, and the ring-electrode coupling. The latter can be selectively tuned in a suitable laboratory setting and has already been examined in various junction setups.
The results analyzed so far are computed for zero temperature. For a more realistic scenario it is also crucial to examine the role of temperature. The temperature dependence enters into the current expression through the Fermi-Dirac distribution functions, $f_{S}$ and $f_{D}$, associated with the source and drain electrodes, respectively. The effect of temperature is presented in Fig. 8, where $I_{c}$-$W$ curves are shown for three distinct temperatures. The result of zero temperature is also superimposed for comparison. Apart from a slight reduction of current with temperature, we find that the overall nature of $I_{c}$-$W$ curve remains unaltered. The reduction of current due to temperature can be understood from the following arguments. At absolute zero temperature, the current is obtained by
@@
integrating the current density profile within the energy window from $E_{F}-eV/2$ to $E_{F}+eV/2$. Within this energy zone, depending on the dominating peaks and dips we get the net current. Here the chance of mutual cancellations is relatively less. On the other hand, when the temperature is finite, we need to consider the full available energy window where the chance of mutual cancellations may slightly increase though it depends on many other factors, especially $W$, and depending on the weight factor ($f_{S}-f_{D}$) we get the resultant current.
@@
$\blacksquare$ Possible ways to detect bias driven circular current: It is quite significant to explore how bias driven circular current can be verified through experimental observations. Detecting such effects in a practical setting presents both challenges and opportunities. Two possible approaches can be considered nt1 for their experimental observation. The first approach involves examining the spectral response of magnetic ions that are positioned either on or near the quantum ring. These ions interact with the local magnetic field induced in their vicinity due to the circular current, and their spectral characteristics can reveal valuable information about the underlying magnetic effects. The second approach focuses on the response of local magnetic moment that develops within the ring itself to an external magnetic field. By studying the behavior of magnetic moment, it is possible to infer details of the induced magnetic field generated by the bias driven circular current. Both methods offer promising avenues for experimentally probing this theoretically studied phenomenon, bridging the gap between theoretical predictions and real experimental observations.
$\blacksquare$ Possible design prescriptions for the junction: Here, it is pertinent to briefly describe how our proposed quantum system can be designed in the laboratory. The system consists of two main components: a nano ring and backbone sites. Several novel strategies are available for fabricating a nano ring, including electron-beam lithography exp1 ; exp2 , ion beam milling exp3 ; exp4 , UV lithography exp5 ; exp6 , annealing techniques exp7 ; exp8 , and more. The backbone sites, which can be considered discrete quantum dots (QDs), are positioned such that each ring site is connected to one backbone site via a tunnel junction. Using suitable gate electrodes, the coupling strength between the ring and the backbone sites can be controlled. Additionally, the backbone site energies can be modulated, either randomly or deterministically, through local gate voltage variations or laser-induced potential fluctuations.
@@ IV Closing Remarks
Using a tight-binding framework to describe the quantum system, all results are derived based on wave-guide theory. The effect of backbone disorder is particularly noteworthy. As the disorder strength increases, the current magnitude initially rises, reaches a maximum, and then decreases, eventually vanishes at very high disorder strengths. These behaviors are thoroughly analyzed with appropriate mathematical results and physical explanations. Additionally, it is established that the effect of disorder remains largely unchanged regardless of the type of disorder. Temperature dependence is also discussed considering a more realistic scenario, showing that significant current can still be obtained over a wide temperature range. At the end, we briefly discuss the detection mechanisms of bias driven circular current and highlight possible strategies for realizing our proposed model in a suitable laboratory.
Finally, we emphasize that if environmental interactions is controlled through suitable laboratory methods, transport behavior can be selectively regulated. This potential for control is highly interesting and important. Moreover, controlling the spin-dependent bias-driven circular current through environmental interaction presents another intriguing phenomenon. Various spin-dependent scattering factors can be considered def1 ; def2 ; def3 ; spi1 ; spi2 , such as spin-moment interactions in magnetic systems, spin-orbit interactions, Zeeman splitting, etc. In each case, distinct non-trivial signatures can be expected. These aspects will be explored in our forthcoming work.
Appendix A Coupled equations involving wave amplitudes at different lattice sites
Referring to Fig. 10, which illustrates a $6$-site ring coupled to source and drain electrodes,
Figure 10: (Color online). Schematic diagram of a ring nanojunction, where a six-site ring is coupled to one-dimensional source and drain electrodes. These electrodes are attached to sites $1$ and $4$ of the ring. Each site of the ring is further connected to a backbone site through a single bond.
here we express all $14$ equations corresponding to the $14$ lattice sites – $6$ associated with the ring sites, $6$ with the backbone sites, and the remaining $2$ with the side-attached electrodes. For the electrodes, these two sites are site number $-1$ of S and site number $1$ of D, as these two sites are directly coupled to the ring sites. The equations are as follows:
$\left(E-\epsilon_{R,1}\right)C_{R,1}$ $=$ $tC_{R,2}+tC_{R,6}+\eta C_{B,1}+\tau_{S}C_{S,-1},$
$\left(E-\epsilon_{R,2}\right)C_{R,2}$ $=$ $tC_{R,1}+tC_{R,3}+\eta C_{B,2},$
$\left(E-\epsilon_{R,3}\right)C_{R,3}$ $=$ $tC_{R,2}+tC_{R,4}+\eta C_{B,3},$
$\left(E-\epsilon_{R,4}\right)C_{R,4}$ $=$ $tC_{R,3}+tC_{R,5}+\eta C_{B,4}+\tau_{D}C_{D,1},$
$\left(E-\epsilon_{R,5}\right)C_{R,5}$ $=$ $tC_{R,4}+tC_{R,6}+\eta C_{B,5},$
$\left(E-\epsilon_{R,6}\right)C_{R,6}$ $=$ $tC_{R,5}+tC_{R,1}+\eta C_{B,6},$
$\left(E-\epsilon_{B,1}\right)C_{B,1}$ $=$ $\eta C_{R,1},$
$\left(E-\epsilon_{B,2}\right)C_{B,2}$ $=$ $\eta C_{R,2},$
$\left(E-\epsilon_{B,3}\right)C_{B,3}$ $=$ $\eta C_{R,3},$
$\left(E-\epsilon_{B,4}\right)C_{B,4}$ $=$ $\eta C_{R,4},$
$\left(E-\epsilon_{B,5}\right)C_{B,5}$ $=$ $\eta C_{R,5},$
$\left(E-\epsilon_{B,6}\right)C_{B,6}$ $=$ $\eta C_{R,6},$
$\left(E-\epsilon_{S,-1}\right)C_{S,-1}$ $=$ $\tau_{S}C_{R,1}+t_{0}C_{S,-2},$
$\left(E-\epsilon_{D,1}\right)C_{D,1}$ $=$ $\tau_{D}C_{R,4}+t_{0}C_{D,2}.$ (8)
Assuming a plane wave incidence of unit amplitude from the source end, the wave amplitude at $m$th site of the source electrode can be expressed as $C_{S,m}=e^{ik(m+1)}+re^{-ik(m+1)}$, while for the outgoing wave in the drain electrode, the wave amplitude is given by $C_{D,m}=\tau e^{ikm}$. Here $r$ and $\tau$ denote the reflection and transmission amplitudes, respectively, and $k$ is the wave vector related to the TB parameters through the energy dispersion relation $E=\epsilon_{0}+2t_{0}\cos(k)$. By solving the above equations, we find all the required coefficients as a function of energy $E$.